# LOCAL: Learning with Orientation Matrix to Infer Causal Structure from Time Series Data

### v2

@@ 4.3. On NetSim Datasets (RQ2)
NGC (Tank et al., 2021) PCMCI (Peters et al., 2013) DYNO (Pamfil et al., 2020) NTS-NO (Sun et al., 2023) TECDI (Li et al., 2023b) LOCAL LOCAL-CNN NGC (Tank et al., 2021) PCMCI (Peters et al., 2013) DYNO (Pamfil et al., 2020) NTS-NO (Sun et al., 2023) TECDI (Li et al., 2023b) LOCAL LOCAL-CNN
@@
Table 5. Performance on CausalTime Data Under AUROC and AUPRC.
@@
GC (Granger, 1969) 0.45$\pm$0.04 0.42$\pm$0.03 0.57$\pm$0.03 0.63$\pm$0.02 0.28$\pm$0.00 0.42$\pm$0.03
SVAR (Sun et al., 2023) 0.62$\pm$0.04 0.63$\pm$0.00 0.71$\pm$0.02 0.79$\pm$0.02 0.58$\pm$0.00 0.68$\pm$0.04
NTS-NO (Sun et al., 2023) 0.57$\pm$0.02 0.63$\pm$0.03 0.71$\pm$0.02 0.71$\pm$0.02 0.58$\pm$0.05 0.46$\pm$0.02
PCMCI (Peters et al., 2013) 0.53$\pm$0.07 0.54$\pm$0.07 0.70$\pm$0.01 0.67$\pm$0.04 0.35$\pm$0.06 0.51$\pm$0.02
Rhino (Gong et al., 2023) 0.67$\pm$0.10 0.63$\pm$0.02 0.65$\pm$0.02 0.76$\pm$0.08 0.38$\pm$0.01 0.49$\pm$0.03
CUTS (Cheng et al., 2023) 0.60$\pm$0.00 0.62$\pm$0.02 0.37$\pm$0.03 0.51$\pm$0.04 0.15$\pm$0.02 0.15$\pm$0.00
CUTS+ (Cheng et al., 2024a) 0.89$\pm$0.02 0.62$\pm$0.07 0.82$\pm$0.02 0.80$\pm$0.08 0.64$\pm$0.12 0.55$\pm$0.13
NGC (Tank et al., 2021) 0.72$\pm$0.01 0.60$\pm$0.01 0.57$\pm$0.01 0.72$\pm$0.01 0.36$\pm$0.05 0.46$\pm$0.01
NGM (Bellot et al., 2022) 0.67$\pm$0.02 0.47$\pm$0.01 0.56$\pm$0.02 0.48$\pm$0.02 0.28$\pm$0.01 0.47$\pm$0.02
LCCM (Brouwer et al., 2021) 0.86$\pm$0.07 0.55$\pm$0.03 0.80$\pm$0.02 0.93$\pm$0.02 0.59$\pm$0.05 0.76$\pm$0.02
eSRU (Khanna and Tan, 2020) 0.83$\pm$0.03 0.60$\pm$0.02 0.76$\pm$0.04 0.72$\pm$0.03 0.49$\pm$0.03 0.74$\pm$0.06
SCGL (Xu et al., 2019) 0.49$\pm$0.05 0.59$\pm$0.06 0.50$\pm$0.02 0.36$\pm$0.03 0.45$\pm$0.03 0.48$\pm$0.02
TCDF (Nauta et al., 2019) 0.41$\pm$0.02 0.50$\pm$0.00 0.63$\pm$0.04 0.65$\pm$0.01 0.36$\pm$0.00 0.55$\pm$0.03
@@
Table 6. Ablation Study on Synthetic Data with $d=\{20,50,100\}$ Nodes, $T=1000$ Timestep, $p$-order =1
@@
To answer RQ2, we conduct the NetSim experiment, which depicts a dataset characterized by high complexity and nonlinear attributes. We show AUROC and AUPRC results in Table 4. Notably, LOCAL demonstrates significant superiority over the baseline method in both linear and nonlinear scenarios, showcasing an average increase of 4.11% in AUPRC (12.33% for LOCAL-CNN) and an average increase of 9.59% in AUROC (12.33% for LOCAL-CNN) while saving 66.63% of time. It is worth mentioning that, due to the ground truth of NetSim not adhering to the low-rank assumption, we slightly increase the embedding dimension to $k=20$ for LOCAL-CNN. This adjustment enables LOCAL-CNN to effectively capture the intricate relationships within the dataset, contributing to its enhanced performance compared to the baseline method.
4.4. On CausalTime Datasets (RQ3)
@@
To answer RQ3, we employ a real-world dataset CausalTime to evaluate the performance of LOCAL. A detailed comparison will be presented in Table 5. We can observe that: (1) Among all the compared methods, LOCAL performs best in the Traffic subset and second-best in Medical, which demonstrates the effectiveness of the proposed method. (2) Our LOCAL achieves the average best performance, which is attributed to the helpful role of explicitly instantaneous and lagged causality information. (3) The linear implementation of LOCAL is more suitable for the CausalTime dataset than the nonlinear version, one possible reason is that the causal relationship between the data appears linear.
4.5. Ablation and Parameter Study (RQ4)
@@
To answer RQ4, we compare LOCAL with three ablation methods. Initially, as presented in Table 6, causal relationships are extracted solely from a $d\times d$ matrix. Subsequently, we enhance the basic low-rank structure by incorporating interventional data and an Asymptotic Causal Mask, leading to substantial performance improvements. Finally, the optimal performance is achieved when all components are integrated into the LOCAL model. This progressive integration of components is illustrated in Figure 5, highlighting the incremental enhancements in performance as each component is added to the model.
Figure 5. Illustration of ablation study on Synthetic data with $d=\{20,50,100\}$ nodes, $T=1000$ timestep, $p$-order =1
@@
Figure 6. The minimum dimension $k$ required for synthetic data with order = 1 to achieve 90% accuracy under different ranks (5%, 10%, 20%, 50%, 60%) for $d=20$ nodes and $T=1000$ time steps.
@@
Figure 6 illustrates that to achieve at least 90% accuracy (using TPR as the evaluation metric), the minimum dimension $k$ increases as the rank-to-$d$ ratio rises, and is approximately proportional to the true rank of the graph structure. This trend may be linked to the inherent data structure: lower matrix ranks often correspond to graphs with significant centralization. As $k$ increases, graph complexity also rises, necessitating a higher dimension to capture the data’s complexity and diversity effectively. This finding aids in selecting an optimal $k$ to maintain accuracy. When the rank is unknown, choosing $k$ to be slightly less than 0.5 times $d$ can balance performance and computational efficiency.
4.6. Embedding Matrices Visualization (RQ5)
@@
To answer RQ5, this section visualizes the learned source embedding and target embedding in a synthetic dataset with $d=5$. We illustrate the ground truth instantaneous causal graph, source embedding $\boldsymbol{E}_{so}(t)$, and target embedding $\boldsymbol{E}_{to}(t)$ in Figure 7. As we can see, the $d\times d$-dimensional causal matrix $\boldsymbol{W}$ is decomposed into $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$, where each row of the source embedding $\boldsymbol{E}_{so}$ corresponds to the variable embedding result in $\boldsymbol{W}$ as parents. Each row of target embedding $\boldsymbol{E}_{to}$ corresponds to the variable embedding result of children. This is exactly why we named it source and target embedding.
Figure 7. Illustration of node source embedding matrix and target embedding matrix in a DBN with $d=5$ nodes.
5. Conclusion
@@
In this work, we introduce an easy-to-implement variant to the dynamic causal discovery family, LOCAL, particularly tailored for capturing nonlinear, high-dimension dynamic systems rapidly. Specifically, it estimates causal structure and connection strength through a quasi-maximum likelihood-based score function, Asymptotic Causal Mask Learning (ACML) and Dynamic Graph Parameter Learning (DGPL) modules respectively. We evaluate our approach on both synthetic and public benchmark datasets, LOCAL has presented advantages over previous state-of-the-art approaches. In the future, we aim to design additional neural network architectures for LOCAL, including attention-based methods.
Acknowledgements.
@@
This work was supported by Beijing Natural Science Foundation under Grant (No.XXXX).

### v3

@@ 3.3 Results on NetSim Datasets
NGC [38] PCMCI [35] DYNO [34] NTS-NO [37] TECDI [26] LOCAL LOCAL-CNN NGC [38] PCMCI [35] DYNO [34] NTS-NO [37] TECDI [26] LOCAL LOCAL-CNN
@@
Table 5: F test results comparing the previous baseline methods (e.g., TECDI) with our proposed methods (LOCAL and LOCAL-CNN). ”$F$” represents the test statistic, while ”$p$” denotes the probability of obtaining the observed results under the null hypothesis. A $p$-value below the significance level (e.g., 0.05) indicates significant differences between the compared algorithms.
Method comparison $F$ $p$
TECDI vs. LOCAL 11.87 0.0015
TECDI vs. LOCAL-CNN 14.53 0.0006
Table 6: Evaluation results of different causal discovery methods on the CausalTime dataset. The performance is assessed using AUROC and AUPRC.
@@
GC [21] 0.45$\pm$0.04 0.42$\pm$0.03 0.57$\pm$0.03 0.63$\pm$0.02 0.28$\pm$0.00 0.42$\pm$0.03
SVAR [37] 0.62$\pm$0.04 0.63$\pm$0.00 0.71$\pm$0.02 0.79$\pm$0.02 0.58$\pm$0.00 0.68$\pm$0.04
NTS-NO [37] 0.57$\pm$0.02 0.63$\pm$0.03 0.71$\pm$0.02 0.71$\pm$0.02 0.58$\pm$0.05 0.46$\pm$0.02
PCMCI [35] 0.53$\pm$0.07 0.54$\pm$0.07 0.70$\pm$0.01 0.67$\pm$0.04 0.35$\pm$0.06 0.51$\pm$0.02
Rhino [20] 0.67$\pm$0.10 0.63$\pm$0.02 0.65$\pm$0.02 0.76$\pm$0.08 0.38$\pm$0.01 0.49$\pm$0.03
CUTS [13] 0.60$\pm$0.00 0.62$\pm$0.02 0.37$\pm$0.03 0.51$\pm$0.04 0.15$\pm$0.02 0.15$\pm$0.00
CUTS+ [11] 0.89$\pm$0.02 0.62$\pm$0.07 0.82$\pm$0.02 0.80$\pm$0.08 0.64$\pm$0.12 0.55$\pm$0.13
NGC [38] 0.72$\pm$0.01 0.60$\pm$0.01 0.57$\pm$0.01 0.72$\pm$0.01 0.36$\pm$0.05 0.46$\pm$0.01
NGM [4] 0.67$\pm$0.02 0.47$\pm$0.01 0.56$\pm$0.02 0.48$\pm$0.02 0.28$\pm$0.01 0.47$\pm$0.02
LCCM [5] 0.86$\pm$0.07 0.55$\pm$0.03 0.80$\pm$0.02 0.93$\pm$0.02 0.59$\pm$0.05 0.76$\pm$0.02
eSRU [23] 0.83$\pm$0.03 0.60$\pm$0.02 0.76$\pm$0.04 0.72$\pm$0.03 0.49$\pm$0.03 0.74$\pm$0.06
SCGL [39] 0.49$\pm$0.05 0.59$\pm$0.06 0.50$\pm$0.02 0.36$\pm$0.03 0.45$\pm$0.03 0.48$\pm$0.02
TCDF [31] 0.41$\pm$0.02 0.50$\pm$0.00 0.63$\pm$0.04 0.65$\pm$0.01 0.36$\pm$0.00 0.55$\pm$0.03
@@
Table 7: Ablation study on the components of LOCAL using a synthetic dataset with $d=\{20,50,100\}$ nodes, $T=1000$ timesteps, and $p$-order = 1. The results demonstrate that each component contributes to improved performance of LOCAL.
@@
We conducted the NetSim experiment, which depicted a dataset characterized by high complexity and nonlinear attributes. We showed AUROC and AUPRC results in Table 4. Notably, LOCAL demonstrated significant superiority over the baseline method in both linear and nonlinear scenarios, showcasing an average increase of 4.11% in AUPRC (12.33% for LOCAL-CNN) and an average increase of 9.59% in AUROC (12.33% for LOCAL-CNN) while saving 66.63% of time. It was worth mentioning that, due to the ground truth of NetSim not adhering to the low-rank assumption, we slightly increased the embedding dimension to $k=20$ for LOCAL-CNN. This adjustment enabled LOCAL-CNN to effectively captured the intricate relationships within the dataset, contributing to its enhanced performance compared to the baseline method.
We presented the F test results in Table 5. The findings indicated that both LOCAL and LOCAL-CNN significantly outperformed TECDI. Specifically, the test statistic ($F$) values of 11.87 and 14.53, with corresponding $p$-values of 0.0015 and 0.0006, provided strong evidence against the null hypothesis of no difference between the methods. As both $p$-values are well below the standard significance threshold (0.05), we concluded that the proposed methods demonstrated superior performance compared to TECDI.
3.4 Results on CausalTime Datasets
@@
We employed a real-world dataset CausalTime to evaluate the performance of LOCAL. A detailed comparison was presented in Table 6. We observed that: (1) Among all the compared methods, LOCAL performed best in the Traffic subset and second-best in Medical, which demonstrated the effectiveness of the proposed method. (2) LOCAL achieved the average best performance, which was attributed to the helpful role of explicitly instantaneous and lagged causality information. (3) The linear implementation of LOCAL was more suitable for the CausalTime dataset than the nonlinear version, one possible reason was that the causal relationship between the data appeared linear.
3.5 Ablation Study
@@
We compared LOCAL with three ablation methods. Initially, as presented in Table 7, causal relationships were extracted solely from a $d\times d$ matrix. Subsequently, we enhanced the basic low-rank structure by incorporating interventional data and an Asymptotic Causal Mask, leading to substantial performance improvements. Finally, the optimal performance was achieved when all components were integrated into the LOCAL model. This progressive integration of components was illustrated in Figure 5, highlighting the incremental enhancements in performance as each component was added to the model.
Figure 5: Ablation study of different components in LOCAL. Each component contribute to the learning of instantaneous dependencies (left) and lagged dependencies (right) on a synthetic dataset with $d=\{20,50,100\}$ nodes, $T=1000$ time steps, and a lag order of $p=1$. The results highlight the impact of removing individual components (DGPL, ACML, and QMLE) on model performance.
@@
Figure 6: Minimum number of latent dimensions $k$ required for LOCAL to achieve 90% accuracy at different rank-to-$d$ ratios. Evaluated on a synthetic dataset with $d=\{20,50,100\}$ nodes, $T=1000$ time steps, and a lag order of $p=1$, the results show that as the rank-to-$d$ ratio increases (5%, 10%, 20%, 50%, 60%), the required latent dimension $k$ also increase.
@@
Figure 6 illustrated that to achieve at least 90% accuracy (using TPR as the evaluation metric), the minimum dimension $k$ increased as the rank-to-$d$ ratio rised, and was approximately proportional to the true rank of the graph structure. This trend might be linked to the inherent data structure: lower matrix ranks often corresponded to graphs with significant centralization. As $k$ increased, graph complexity also rised, necessitating a higher dimension to capture the data’s complexity and diversity effectively. This finding aided in selecting an optimal $k$ to maintain accuracy. When the rank was unknown, choosing $k$ to be slightly less than 0.5 times $d$ could balance performance and computational efficiency.
3.6 Discussion: What do source embedding and target embedding learn respectively?
@@
This section visualized the learned source embedding and target embedding in a synthetic dataset with $d=5$. We illustrated the ground truth instantaneous causal graph, source embedding $\boldsymbol{E}_{so}(t)$, and target embedding $\boldsymbol{E}_{to}(t)$ in Figure 7. As we could see, the $d\times d$-dimensional causal matrix $\boldsymbol{W}$ was decomposed into $\boldsymbol{E}_{so}$ and $\boldsymbol{E}_{to}$, where each row of the source embedding $\boldsymbol{E}_{so}$ corresponded to the variable embedding result in $\boldsymbol{W}$ as parents. Each row of target embedding $\boldsymbol{E}_{to}$ corresponded to the variable embedding result of children. This was exactly why we named it source and target embedding.
Figure 7: Illustration of the node source embedding matrix ($E_{so}$) and target embedding matrix ($E_{to}$) in a dynamic Bayesian network (DBN) with $d=5$ nodes. The leftmost graph represent the ground truth structure, while the matrices visualize the learned embeddings.
4 Conclusion
@@
In this work, we introduce an easy-to-implement variant to the dynamic causal discovery family, LOCAL, particularly tailored for capturing nonlinear, high-dimension dynamic systems rapidly. Specifically, it estimates causal structure and connection strength through a quasi-maximum likelihood-based score function, Asymptotic Causal Mask Learning (ACML) and Dynamic Graph Parameter Learning (DGPL) modules respectively. We evaluate our approach on both synthetic and public benchmark datasets, LOCAL has presented advantages over previous state-of-the-art approaches. In summary, LOCAL offers an efficient and flexible solution for dynamic causal discovery in time series, demonstrating strong potential for enhancing both efficiency and performance.
5 Limitation
While LOCAL demonstrates strong performance across various tasks and datasets, there remain opportunities for further exploration. The DGPL module leverages low-rank matrix factorization to reduce the number of parameters. However, if the true causal structure exhibits a high rank (e.g., densely connected networks), this decomposition may introduce biases in estimating causal strength. Additionally, while our evaluation provides a solid foundation, further research is needed to explore the performance of LOCAL on datasets with a larger number of nodes and longer lagged dependencies. In future work, deep time series modeling, adaptive rank optimization, and non-stationary analysis can be integrated to better capture the causal dynamics of complex real-world systems.
Acknowledgements
@@
This work was supported by the National Key Research and Development Program of China under Grant (No.2024YFB3312200), and by the Beijing Natural Science Foundation under Grant (No.L231005).
Ethical Statement
1. Articles do not rely on clinical trials.
2. The authors declare that they have no conflict of interest.
3. All submitted manuscripts containing research which does not involves human participants and/or animal experimentation.